health insurance policy pays 70 percent of physical therapy costs after a $200 deductible. In contrast, an HMO charges $17 per visit for physical therapy. How much would a person save with the HMO if he or she had 10 physical therapy sessions costing $50 each? (Do not round intermediate calculations.)
Answer #1

Cost with health insurance=200+(1-70%)*(10*50-200)=290
Cost with HMO: 10 sessions=10*17=170
Savings of 290-170=120 with HMO